OpenStudy (anonymous):

What is the sum of the geometric sequence 1, 4, 16, if there are 8 terms? (1 points) 21,845 42,372 55,987 87,381

OpenStudy (anonymous):

it is each previous number times 4 1 4 16 64 256 1024 4096 16384 add these up and you will have your answer

OpenStudy (anonymous):

It Gives Me 21845

OpenStudy (atlas):

S = 1 + 4 + 16 +........8 terms or, S = 4^0 + 4^1 + 4^2 + ............+4^7 .......(i) or,4 S = 4^1 + 4^2 +.............+4^7 + 4^8 .......(II) Subtract i from ii: 3S = 4^8 - 4^0 or S = (4^8 -1)/3

OpenStudy (atlas):

21845 is the right answer

OpenStudy (anonymous):

Can U Check One For Me , I Did It What is the sum of the geometric sequence 1, -4, 16, if there are 7 terms? (1 points) -5,461 5,461 -3,277 3,277 I Got 3277

OpenStudy (anonymous):

@atlas @cbobo

OpenStudy (atlas):

The general formula for a geometric sequence is S = a(r^n -1)/(r-1) where a is the first term and r the common ratio

OpenStudy (anonymous):

so its -3277 , @atlas